OpenStudy (anonymous):

5-square root of 6 over 9.

OpenStudy (anonymous):

\[\frac{5 - \sqrt{6}}{9}\]??? You could add parens for clarity.

OpenStudy (anonymous):

simplify

OpenStudy (anonymous):

I was asking if that was the question, or if it was something else like \[5-\frac{\sqrt{6}}{9}\] because your wording was ambiguous.

OpenStudy (anonymous):

what u first typed

OpenStudy (anonymous):

are you supposed to calculate square root of 6 = 2.45 and get the final answer of 0.2834, or are you supposed to find the simplest form in terms of a fraction with radicals?

OpenStudy (anonymous):

rationalize numerator

OpenStudy (anonymous):

so multiply by\[\frac{5+\sqrt{6}}{5+\sqrt{6}}\]

OpenStudy (anonymous):

and get?

OpenStudy (anonymous):

numerator: 19 denominator: 45 + 9sqrt(6)